OpenStudy (anonymous):

For some positive constant C, a patient's temperature change T, due to a dose, D, of a drug is given by T= (C/2 - D/3)D^2. What dosage maximizes the temperature change? The sensitivity of the body to the drug is defined as dT/dD. What dosage maximizes sensitivity? For some positive constant C, a patient's temperature change T, due to a dose, D, of a drug is given by T= (C/2 - D/3)D^2. What dosage maximizes the temperature change? The sensitivity of the body to the drug is defined as dT/dD. What dosage maximizes sensitivity? @Mathematics

OpenStudy (dumbcow):

to maximize T , set dT/dD = 0 to maximize sensitivity, set dT^2/dD = 0 **2nd derivative** dT/dD = CD -D^2 dT^2/dD = C -2D CD -D^2 = 0 D(C-D) = 0 Assume D>0 C-D = 0 D=C --> when dosage equals C, T is maximized C -2D = 0 D = C/2 --> when dosage equals C/2, sensitivity is maximized